# Covering Complete Geometric Graphs by Monotone Paths

### v1

@@ Abstract
Given a set $A$ of $n$ points (vertices) in general position in the plane, the complete geometric graph $K_{n}[A]$ consists of all (${n\choose 2}$ segments (edges) between the elements of $A$. It is known that the edge set of every complete geometric graph on $n$ vertices can be partitioned into $O(n^{3/2})$ noncrossing paths (or matchings). We strengthen this result under various additional assumptions on the point set. In particular, we prove that for a set $A$ of $n$ randomly selected points, uniformly distributed in $[0,1]^{2}$, with probability tending to $1$ as $n\rightarrow\infty$, the edge set of $K_{n}[A]$ can be covered by $O(n^{4/3}t(n))$ noncrossing paths (or matchings), where $t(n)$ is any function tending to $\infty$. On the other hand, we construct $n$-element point sets such that covering the edge set of $K_{n}(A)$ requires a quadratic number of monotone paths.
@@ 1 Introduction
A set of points in the plane is said to be (i) in general position if no $3$ points are collinear; and (ii) in convex position if every point in the set is an extreme point of the convex hull. Given $n$ points in general position in the plane, the graph obtained by connecting certain point-pairs by straight-line segments is called a geometric graph $G$. If no two segments (edges) of $G$ cross each other, then $G$ is said to be noncrossing or a plane graph.
In $2005$, Araujo, Dumitrescu, Hurtado, Noy, and Urrutia [3] asked the following question: Can the edge set of every complete geometric graph on $n$ vertices be partitioned into a small number of noncrossing matchings? In particular, they asked:
@@
It is easy to verify that when $n$ is even (resp., odd), every complete geometric graph of $n$ vertices in convex position can be decomposed into $n-1$ (resp., $n$) perfect plane matchings. On the other hand, the best known upper bound for point sets in general position is only $O(n^{3/2})$ [3], which seems to be far off.
In this paper we study a variant of the above question, namely the problem of covering complete geometric graphs by monotone paths, a special type of plane paths. A path (or matching) is called monotone if there is a straight line $\ell$ such that no line orthogonal to $\ell$ intersects it in more than one point. How many paths does one need in this case? Note that every path naturally decomposes into two matchings.
Monotone paths are ubiquitous and have been studied extensively over the years, particularly in the field of optimization [6]. On the other hand, noncrossing paths and matchings have been an attraction for geometers [8, 9, 10]. Moreover, every monotone path is noncrossing.
Before stating our results, we clarify the terminology.
Definitions and notations.
For a triple of points $a,b,c$, let $\Delta{abc}$ denote the triangle with vertices $a,b,c$. All point sets appearing in this paper are assumed to be in general position, and the logarithms are in base $2$.
A polygonal path $\xi=(v_{1},v_{2},\ldots,v_{t})$ in $\mathbb{R}^{2}$ is monotone in direction $\mathbf{u}\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$ (or is $\mathbf{u}$-monotone) if every directed edge of $\xi$ has a positive inner product with $\mathbf{u}$, that is, $\langle\overrightarrow{v_{i}v_{i+1}},\mathbf{u}\rangle>0$ for $i=1,\ldots,t-1$; here $\mathbf{0}$ is the origin. A path $\xi=(v_{1},v_{2},\ldots,v_{t})$ is monotone if it is monotone in some direction $\mathbf{u}\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$. Obviously, any monotone path is noncrossing.
@@
Our results.
Theorem 1.1.
Let $A$ be a set of $n$ random points uniformly distributed in $[0,1]^{2}$, and let $n\rightarrow\infty$. Let $t(n)$ be an arbitrarily slowly growing function of $n$, with $\lim_{n\to\infty}t(n)=\infty$.
Then, with probability tending to $1$, the edge set of $K_{n}[A]$ can be covered by at most $O(n^{4/3}t(n))$ noncrossing paths, and, hence, by $O(n^{4/3}t(n))$ noncrossing matchings.
This is better than the $O(n^{3/2})$ bound of Araujo et al. [3], which holds for any point set in general position.
Note that a covering of the edge set of $K_{n}[A]$ by $O(n^{4/3}t(n))$ noncrossing matchings yields a decomposition (i.e., partition) of $K_{n}[A]$ into $O(n^{4/3}t(n))$ noncrossing matchings (via deletion of duplicate edges). However, the same is not true for coverings by paths.
@@
Let $A$ be an $\alpha$-dense point set in general position in the plane.
The edge set of $K_{n}[A]$ can be covered by $O(n^{3/2})$ monotone paths, and such a covering can be computed in $O(n^{2})$ time. Consequently, the edge set of $K_{n}[A]$ can be decomposed into $O(n^{3/2})$ monotone matchings.
From the opposite direction, we suspect that covering the edge set of the complete geometric graph of any dense point set requires a superlinear number of monotone paths.
@@
For some (possibly, for any) dense $n$-element point set $A$ in general position in the plane, covering the edge set of $K_{n}[A]$ requires a superlinear number of monotone paths.
@@
Let $A$ be a set of $n$ points in general position in the plane.
Then the edge set of the complete geometric graph $K_{n}[A]$ can be covered by $n^{2}/6+O(n)$ monotone paths. On the other hand, there exist $n$-element point sets that require at least $n^{2}/15$ monotone paths.
The rest of this paper is organized as follows. In Sections 2, 3, and 4, we prove Theorems 2, 3, and 5, respectively. The last section contains some concluding remarks and open problems.
@@
Let $A=a_{1},\ldots,a_{n}$ be a sequence of $n$ points, randomly and uniformly distributed in $U=[0,1]^{2}$. Write $m={n\choose 2}$ and let $e_{1},\ldots,e_{m}$ be the edges of $K_{n}[A]$ listed in lexicographic order. Consider the following procedure for covering the edge set of $K_{n}[A]$ by monotone paths.
Let $\theta=cn^{-4/3}$ be a small angle, for a suitable small constant $c>0$, so that $N=\pi/\theta$ is an integer. Observe that sorting with respect to $\mathbf{v}$ and $-\mathbf{v}$ are equivalent. Let $\mathcal{V}$ be a set of $N$ unit vectors uniformly distributed in $[0,\pi)$ so that $(1,0)\in\mathcal{V}$. Observe that every monotone path naturally decomposes into two noncrossing matchings. For each vector $\mathbf{v}\in\mathcal{V}$, compute the $\mathbf{v}$-monotone spanning path $\xi$ induced by $A$ and output the two $\mathbf{v}$-monotone matchings whose union is $\xi$. Consequently, we have obtained $O(n^{4/3})$ monotone paths (or matchings). This forms Phase I of the covering.
We show that the expected number of edges that do not appear in these $O(n^{4/3})$ matchings is $O(n^{5/3})$. Further decompose the geometric graph $G=(V,E)$ on this edge set into $O(n^{4/3})$ noncrossing matchings (or paths), as outlined below. This forms Phase II of the covering. Consequently, the expected number of noncrossing matchings generated by the above procedure is $O(n^{4/3})$ in each of the two phases, hence $O(n^{4/3})$ overall.
Analysis.
We start with Phase I. Let $a,b\in A$ be any pair of points. Instead of working with monotone paths in many directions, we may assume for convenience that $ab$ is a vertical segment, and that the unit square has been rotated by some angle, correspondingly. We denote by $U_{\alpha}$ the square rotated by $\alpha$; we have ${\rm Area}(U_{\alpha})=1$ for any $\alpha$. Consider the geometric locus $R(a,b)$ of the points $p\in U_{\alpha}$ on lines making an angle of at most $\theta$ with $ab$ and which intersect $ab$, see Fig. 1.
Figure 1: Left: four lines making an angle of $\theta$ with $ab$ through $a$ and $b$. Center: the geometric locus (shaded) associated with a vertical segment. Right: a bounding rectangle for the locus associated with a long vertical segment.
@@
Consider four lines incident to the $a$ and $b$ respectively, and making a (clockwise or counterclockwise) angle of $\theta$ with $ab$. The boundary of the locus is made by these four lines and two polygonal arcs on the boundary of the square. The locus is easily seen to be contained in an axis-parallel rectangle of width $2\sqrt{2}\tan\theta$ and height $\sqrt{2}$. As such, its area is bounded from above by $4\tan\theta$. Excluding the two isosceles triangles based on the left and right vertical sides yields an improved area bound of $\frac{3}{4}\cdot 4\tan\theta=3\tan\theta\leq 4\theta$ (here we assume that $n$ is sufficiently large).
Observe that if the region $R(a,b)$ is empty of points in $A\setminus\{a,b\}$, then the pair $a,b$ generates an edge in any $\mathbf{v}$-monotone spanning path induced by $A$, such that $\mathbf{v}\in\mathcal{V}$ makes an angle at most $\theta$ with the direction orthogonal to $ab$. In this case, $ab$ appears in part I of the decomposition. We quantify the probability of this event.
We say that pair $a,b$ is good if $ab$ is an edge of a $\mathbf{v}$-monotone spanning path considered in Phase I, with $\mathbf{v}$ making an angle at most $\theta$ with a line orthogonal to $ab$; and it is bad otherwise. By the principle of deferred decisions, see, e.g., [12], we may assume that $a,b$ are the first two points in the sequence of $n$ random draws when considering this particular $\mathbf{v}$. By using the independence of trials and Bernoulli’s inequality, we obtain:
${\rm Prob}\left(a,b\text{ is good}\right)$ $\geq\left(\frac{1-{\rm Area}(R(a,b)}{{\rm Area}(U_{\alpha})}\right)^{n-2}$
$>(1-{\rm Area}(R(a,b))^{n}\geq(1-4\theta)^{n}\geq 1-4n\theta$
$\geq 1-4cn^{-1/3}.$
It follows that
${\rm Prob}\left(a,b\text{ is bad}\right)\leq 4cn^{-1/3}.$ (1)
Recall that $e_{1},\ldots,e_{m}$ are the edges of $K_{n}[A]$ in a fixed (lexicographic) order. Let $Y_{i}$, $i=1,\ldots,m$ be the indicator random variable of the event that $e_{i}$ is a bad edge and let $Y=\sum_{i=1}^{m}Y_{i}$. By (1), the expected number of edges not appearing in monotone paths generated in Phase I is bounded from above as follows:
${\rm E}[Y]\leq m\cdot 4cn^{-1/3}\leq 2cn^{5/3}=:\mu.$
We use Markov’s inequality, see, e.g., [12, Thm. 3.1], to show that the bound we obtain on the number of monotone paths holds with probability tending to $1$ as $n\to\infty$.
${\rm Prob}\left(Y\geq\mu\cdot t(n)\right)\leq(t(n))^{-1}\to 0.$ (2)
We continue with Phase II. Let $G=(V,E)$, with $|E|=Y$, be the random geometric graph of bad edges surviving Phase I. For decomposing the edge set of $G$ into noncrossing matchings (or paths), we use a result of Tóth [19], according to which any geometric graph on $n$ points and with no $k+1$ disjoint edges has at most $2^{9}k^{2}n=O(k^{2}n)$ edges. Moreover, if the graph has more than $2^{9}k^{2}n$ edges, one can find $k+1$ pairwise disjoint edges, or a noncrossing path of $12k+$ 1 edges.
In our case, solving for $k$ in $n^{5/3}=nk^{2}$ yields $k=n^{1/3}$. Start with the graph $G$ on the $n$ points, and repeatedly remove a large set of pairwise disjoint edges, until the graph becomes empty. The process consists of at most $\lceil\log{(2cn^{5/3})}\rceil$ steps, $i=2,\ldots,\lceil\log{(2cn^{5/3})}\rceil+1$. In step $i$, the current graph has $m_{i}$ edges, where
$\frac{2cn^{5/3}}{2^{i}}<m_{i}\leq\frac{2cn^{5/3}}{2^{i-1}}.$
In step $i$, repeatedly find and remove $k_{i}+1$ disjoint edges (a matching), or a path of $22k_{i}+$ 2 edges, where $k_{i}\sim n^{1/3}2^{-i/2}$, until $m_{i}$ fails to satisfy the above inequality, and the process continues with step $i+1$. The number of matchings (or paths) removed in step $i$ is about
$\frac{n^{5/3}}{2^{i}}\cdot\frac{2^{i/2}}{n^{1/3}}=n^{4/3}\cdot 2^{-i/2},$
and so the total number of matchings (or paths) is at most
$n^{4/3}\sum_{i=2}^{\infty}2^{-i/2}=O(n^{4/3}).$
A final note: the paths (or matchings) that are generated by the procedure in [19] are not necessarily monotone. ∎
@@
Consider an $\alpha$-dense point set $A$. By scaling, we can assume that $A$ is contained in a unit square $U=[0,1]^{2}$, and the distance between any two points of $A$ is at least $\sqrt{2}\cdot\alpha n^{-1/2}$.
The covering procedure is similar to Phase I in the proof of Theorem 1.1, but with a twist. Let $\varphi=cn^{-1/2}$ be a small angle, for a suitable small constant $c>0$, so that $N=2\pi/\varphi$ is an integer. Let $\mathcal{V}$ be a set of $N$ unit vectors uniformly distributed around the origin so that $(1,0)\in\mathcal{V}$. For each vector $\mathbf{v}\in\mathcal{V}$, compute the $\mathbf{v}$-monotone spanning path $\xi_{v}$ induced by $A$. Let the path be $a_{1}a_{2}\ldots a_{n}$ in the order of monotonicity. Set
$\beta=\frac{8c+4\sqrt{2}\cdot\alpha}{\pi\alpha^{2}},$
and for each $1\leq j\leq i\leq\beta\sqrt{n}$, add the path $a_{j}a_{i+j}a_{2i+j}\ldots$ to the set $\mathcal{P}$ of covering paths. Note that for each vector $\mathbf{v}\in\mathcal{V}$, there are at most $1+2+\cdots+\beta\sqrt{n}<\beta^{2}n$ paths to be added to $\mathcal{P}$, all of which are monotone and therefore $\mathcal{P}$ contains at most $\frac{2\pi}{c}\sqrt{n}\cdot\beta^{2}n=\Theta(n^{3/2})$ monotone paths.
We claim that the paths in $\mathcal{P}$ cover the complete geometric graph on $A$. That is, for any pair of points $a,b\in A$, we show that the segment $ab$ appears in at least one path in $\mathcal{P}$. We construct the region $R(a,b)$ analogous to the one in the proof of Theorem 1.1 but with respect to $\varphi$ instead. By Lemma 2.1, its area is at most $4\varphi$. Assume that $R(a,b)$ contains $k$ points of $A$. The disks of radius $r=\frac{\sqrt{2}}{2}\cdot\alpha n^{-1/2}$ centered at these $k$ points are disjoint and they cannot exceed the region $R(a,b)$ extended by a margin of $r$, whose area is then at most $4(\varphi+r)$. The total area of the $k$ disks is at most $4(\varphi+r)$, and therefore
@@
$k\pi\frac{\alpha^{2}}{2n}$ $\leq 4\left(\frac{c}{\sqrt{n}}+\frac{\frac{\sqrt{2}}{2}\cdot\alpha}{\sqrt{n}}\right),\text{ or }$
$k$ $\leq\frac{8c+4\sqrt{2}\alpha}{\pi\alpha^{2}}\cdot\sqrt{n}=\beta\sqrt{n}.$
@@
Lower bound. Assume that $n$ is divisible by $3$, i.e., $n=3k$. For convenience we denote our $n$-element point set by $P$. Partition the $n$ points into three groups $A,B,C$, of the same size, close to the vertices of a unit equilateral triangle, and each group consists of roughly equidistant points, but the distances are different; see Fig. 2. Let the diameters of the 3 groups of $k$ vertices be ${\rm diam}(C)\ll{\rm diam}(B)\ll{\rm diam}(A)\ll 1$.
The number of (undirected) inter-group edges, i.e., edges in
@@
is $33k^{2}=n^{2}/$ 3. Let $\mathcal{P}$ be a covering of $K_{n}[A]$ by monotone paths. We claim that every path $\xi\in\mathcal{P}$ contains at most five edges in $E_{0}$. Consequently, covering all the edges of the tripartite graph on $P=A\cup B\cup C$ requires at least $153k^{2}/5=n^{2}/$ 15 monotone paths.
@@
Upper bound. Since every 2-edge path is monotone, one can decompose the edge set of $K_{n}[A]$ into $\lceil\frac{1}{2}{n\choose 2}\rceil$ 2-edge paths. We can improve upon this bound by using $3$-edge paths. Since not every $3$-edge path is monotone, we need special $3$-edge paths that are monotone. A $3$-edge path is a zig-zag path if its two endpoints lie in opposite open halfplanes determined by the middle edge of the path. It is easy to see that any $3$-edge zig-zag path is monotone.
@@
Let $n$ be a large enough positive integer. Then the edge set of $K_{n}$ can be decomposed into subgraphs of the form $K_{6}$, if and only if $n$ satisfies two divisibility conditions, namely $15|n-$ 1 and $\binom{6}{2}|\binom{n}{2}$.
@@
Figure 3: Covering the edge set of the complete graph on six points by $3$-edge zig-zag paths.
Our approach is packing as many $K_{6}$ as possible into $K_{n}[A]$, then decomposing each $K_{6}$ into five $3$-edge zig-zag paths, and considering each of the remaining edges as a single path. This way we can guarantee a decomposition of the edge set of $K_{n}[A]$ into $\binom{n}{2}/\binom{6}{2}\times 5+O(n)=n^{2}/6+O(n)$ monotone paths.
It suffices to show that for a $6$-point set $A$, the edge set of $K_{6}[A]$ can be decomposed into five $3$-edge zig-zag paths. It is known that there are $16$ order types for such sets, see, e.g., [17], and it suffices to provide zig-zag path decompositions for each type, as we do in Fig. 3. ∎
A similar bipartite version of the lower bound construction with $n=2k$ points and $|E_{0}|=k^{2}$ (undirected) inter-group edges only yields a lower bound of $n^{2}/16$: If $\mathcal{P}$ is a covering of $K_{n}[A]$ by monotone paths, it has the property that that every path $\xi\in\mathcal{P}$ contains at most four edges in $E_{0}$. Consequently, covering all the edges of the bipartite graph requires at least $k^{2}/4=n^{2}/16$ monotone paths.

### v2

@@ Abstract
Given a set $A$ of $n$ points (vertices) in general position in the plane, the complete geometric graph $K_{n}[A]$ consists of all $\binom{n}{2}$ segments (edges) between the elements of $A$. It is known that the edge set of every complete geometric graph on $n$ vertices can be partitioned into $O(n^{3/2})$ crossing-free paths (or matchings). We strengthen this result under various additional assumptions on the point set. In particular, we prove that for a set $A$ of $n$ randomly selected points, uniformly distributed in $0,1]^{2}$, with probability tending to $1$ as $n\rightarrow\infty$, the edge set of $K_{n}[A]$ can be covered by $O(n\log n)$ crossing-free paths and by $O(n\sqrt{\log n})$ crossing-free matchings. On the other hand, we construct $n$-element point sets such that covering the edge set of $K_{n}[A]$ requires a quadratic number of monotone paths.
@@ 1 Introduction
A set of points in the plane is said to be (i) in general position if no $3$ points are collinear; and (ii) in convex position if every point in the set is an extreme point of the convex hull. Given $n$ points in general position in the plane, the graph obtained by connecting certain point-pairs by straight-line segments is called a geometric graph $G$. If no two segments (edges) of $G$ cross each other, then $G$ is said to be crossing-free or a plane graph.
In $2005$, Araujo, Dumitrescu, Hurtado, Noy, and Urrutia [3] asked the following question: Can the edge set of every complete geometric graph on $n$ vertices be partitioned into a small number of crossing-free matchings? In particular, they asked:
@@
It is easy to verify that when $n$ is even (respectively odd), every complete geometric graph of $n$ vertices in convex position can be decomposed into n1n-1 (respectively $n$) perfect plane matchings. On the other hand, the best known upper bound for point sets in general position is only $O(n^{3/2})$ [3], which seems to be far off.
Theorem 1.1.
Let $A$ be a set of $n$ random points uniformly distributed in $0,1]^{2}$, and let $n\rightarrow\infty$. Then, with probability tending to $1$, the edge set of $K_{n}[A]$ can be covered by at most $O(n\log n)$ crossing-free paths, and by $O(n\sqrt{\log n})$ crossing-free matchings.
This is better than the $O(n^{3/2})$ bound in [3], which holds for every point set in general position. The proof in fact gives the stronger result that there is a covering by $O(n\log n)$ monotone paths.
Definition 1.2.
A polygonal path $\xi=(v_{1},v_{2},\ldots,v_{t})$ in $\mathbb{R}^{2}$ is monotone in direction $\mathbf{u}\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$ (or is $\mathbf{u}$-monotone) if every directed edge of $\xi$ has a positive inner product with $\mathbf{u}$, that is, $\langle\overrightarrow{v_{i}v_{i+1}},\mathbf{u}\rangle>0$ for $i=1,\ldots,t-1$; here $\mathbf{0}$ is the origin. A path $\xi=(v_{1},v_{2},\ldots,v_{t})$ is monotone if it is monotone in some direction $\mathbf{u}\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$. Obviously, every monotone path is crossing-free.
@@
Let $A$ be an $\alpha$-dense point set in general position in the plane. Then the edge set of $K_{n}[A]$ can be covered by $O(n^{3/2})$ monotone paths, and such a covering can be computed in $O(n^{2})$ time. Consequently, the edge set of $K_{n}[A]$ can be decomposed into $O(n^{3/2})$ monotone matchings.
From the opposite direction, we suspect that covering the edge set of the complete geometric graph of every dense point set requires a superlinear number of monotone paths.
@@
For some (possibly, for every) dense $n$-element point set $A$ in general position in the plane, covering the edge set of $K_{n}[A]$ requires a superlinear number of monotone paths.
@@
Let $A$ be a set of $n$ points in general position in the plane. Then the edge set of the complete geometric graph $K_{n}[A]$ can be covered by $n^{2}/6+O(n)$ monotone paths. On the other hand, there exist $n$-element point sets that require at least $n^{2}/15$ monotone paths.
Monotone paths are ubiquitous and have been studied extensively over the years, particularly in the field of optimization [6]. On the other hand, crossing-free paths and matchings have been an attraction for geometers [8, 9, 10]. Note that every monotone path is crossing-free. Clearly, by taking the odd, resp. even numbered edges, every monotone path decomposes into two monotone matchings.
The rest of this paper is organized as follows. In Sections 2, 3, and 4, we prove Theorems 2, 3, and 5, respectively. The last section contains some concluding remarks and open problems. All point sets appearing in this paper are assumed to be in general position, and the logarithms are in base $2$.
@@
Let $A=a_{1},\ldots,a_{n}$ be a random sequence of $n$ points, independently and uniformly distributed in $U=[0,1]^{2}$. For any unit vector $\mathbf{v}\in{\mathbb{R}}^{2}$, we can sort $A$ according to the projections $\langle a_{i},\mathbf{v}\rangle$. Observe that sorting with respect to $\mathbf{v}$ and $\mathbf{v}$ are equivalent. Let $\theta=\alpha(\log n)/n$ be a small angle, where $\alpha>0$ is a suitable small constant, so that $N=\pi/\theta$ is an integer. Let $\mathcal{V}$ be a set of $N$ unit vectors evenly spaced in $0,\pi)$ so that $(1,0)\in\mathcal{V}$.
Consider the following procedure for covering the edge set of $K_{n}[A]$ by (i) monotone paths and (ii) crossing-free matchings. For each vector $\mathbf{v}\in\mathcal{V}$, sort the elements of $A$ according to their projection on $\mathbf{v}$, and label them in increasing order as $a_{1}^{\mathbf{v}},\dots,a_{n}^{\mathbf{v}}$. Now proceed as follows:
  • $\bullet$
    To produce a collection of monotone paths: for each $\mathbf{v}\in\mathcal{V}$, and $1\leq i\leq j\leq\beta\log n$, we consider the sequence $(a_{k}^{\mathbf{v}}:k\equiv i\bmod j)$. This defines a monotone path for each choice of $i$, $j$ and $\mathbf{v}$.
  • $\bullet$
    To produce a collection of non-crossing matchings: we partition $A$ into $t:=\lceil n/(\beta\log n)\rceil$ intervals $B_{1},B_{2}\dots B_{t}$ of size at most $\lceil\beta\log n\rceil$. For each set of the form $B_{i}\cup B_{i+1}$, by the $O(n^{3/2})$ aforementioned bound in [3], we can cover the complete graph on $B_{i}\cup B_{i+1}$ with at most $(2\lceil\beta\log n\rceil)^{3/2}$ matchings. We use these to produce a total of $2(2\lceil\beta\log n\rceil)^{3/2}$ matchings: first pair up the matchings from $B_{1}\cup B_{2},B_{3}\cup B_{4},\dots$, and then pair up the matchings from $B_{2}\cup B_{3},B_{4}\cup B_{5},\dots$.
For each choice of $\mathbf{v}$, the first bullet produces a set of at most $(\beta^{2}/2)\log^{2}n$ paths; and the second bullet produces a set of at most $2(2\lceil\beta\log n\rceil)^{3/2}$ matchings. Since there are $N=n\pi/\alpha\log n$ choices of $\mathbf{v}$, we use a total of at most $O(n\log n)$ paths and at most $O(n\sqrt{\log n})$ matchings.
All that remains is to check that the paths and matchings we have produced do in fact cover all pairs of points from $A$. Note that, for each of the orderings we define, we have covered all pairs that are at most $\beta\log n$ apart in the ordering. Thus it is enough to show that, for every pair of points $a,b\in A$, there is some $\mathbf{v}$ such that $a$ and $b$ are at most $\beta\log n$ apart in the corresponding ordering of $A$.
We first note a geometric fact. Let $a,b\in A$ be any two (distinct) points in $U=[0,1]^{2}$. We rotate the picture, and assume for convenience that $ab$ is a vertical segment, and that the unit square has been rotated by some angle. Let $U_{\gamma}$ denote the square rotated by $\gamma$; we have ${\rm Area}(U_{\gamma})=1$ for every $\gamma$. Consider the geometric locus $R(a,b)$ of the points $p\in U_{\gamma}$ on lines making an angle of at most $\theta$ with $ab$ and which intersect $ab$, see Fig. 1.
Figure 1: Left: four lines making an angle of $\theta$ with abab through $a$ and $b$. Center: the geometric locus (shaded) associated with a vertical segment. Right: a bounding rectangle for the locus associated with a long vertical segment.
@@
Consider four lines incident to $a$ and $b$ respectively, and making (clockwise or counterclockwise) angles of $\theta$ with $ab$. The boundary of the locus is made by these four lines and two polygonal arcs on the boundary of the square. The locus is easily seen to be contained in an axis-parallel rectangle of width $2\sqrt{2}\tan\theta$ and height $\sqrt{2}$. As such, its area is bounded from above by $4\tan\theta$. Excluding the two isosceles triangles based on the left and right vertical sides yields an improved area bound of $\frac{3}{4}\cdot 4\tan\theta=3\tan\theta\leq 4\theta$ (here we assume that $n$ is sufficiently large).
We are now ready to complete the proof. We reveal two points $a,b\in A$ (and note that $a\neq b$ with probability 1). Choose a vector $\mathbf{v}\in\mathcal{V}$ that makes an angle at most $\theta$ with the direction orthogonal to $ab$. Then any point of $A$ that comes between $a$ and $b$ in the ordering generated by $\mathbf{v}$ must lie in the region $R(a,b)$, and the lemma shows that, for any given point, the probability that the point lies in this region is at most $4\theta$. The probability that at least $\beta\log n$ points of $A$ lie in $R(a,b)$ is therefore at most
$\binom{n}{\lceil\beta\log n\rceil}(4\theta)^{\lceil\beta\log n\rceil}\leq\left(\frac{en}{\lceil\beta\log n\rceil}\right)^{\lceil\beta\log n\rceil}(4\theta)^{\lceil\beta\log n\rceil}\leq\left(\frac{4\alpha e}{\beta}\right)^{\beta\log n},$
where we used the estimate $\binom{n}{k}\leq(en/k)^{k}$. This probability is $o(1/n^{2})$, provided we have chosen $\alpha$ and $\beta$ sensibly. There are $O(n^{2})$ pairs: so, by the union bound, with high probability, for every pair $a,b\in A$ there is a direction in which $a$ and $b$ are separated by at most $\beta\log n$ other points. This completes the proof. ∎
@@
Consider an $\alpha$-dense point set $A$. By scaling, we can assume that $A$ is contained in the unit square $U=[0,1]^{2}$, and the distance between any two points of $A$ is at least $\sqrt{2}\cdot\alpha^{-1}n^{-1/2}$.
The covering procedure is similar to Phase I in the proof of Theorem 1.1, but with a twist. Let $\varphi=cn^{-1/2}$ be a small angle, for a suitable small constant $c>0$, so that $N=2\pi/\varphi$ is an integer. Let $\mathcal{V}$ be a set of $N$ unit vectors evenly spaced around the origin so that $(1,0)\in\mathcal{V}$. For each vector $\mathbf{v}\in\mathcal{V}$, compute the $\mathbf{v}$-monotone spanning path $\xi_{v}$ induced by $A$. Let the path be $a_{1}a_{2}\ldots a_{n}$ in the order of monotonicity. Set
$\beta=\frac{8\alpha^{2}c+4\sqrt{2}\cdot\alpha}{\pi},$
and for each $1\leq j\leq i\leq\beta\sqrt{n}$, add the path $a_{j}a_{i+j}a_{2i+j}\ldots$ to the set $\mathcal{P}$ of covering paths. Note that for each vector $\mathbf{v}\in\mathcal{V}$, there are at most $1+2+\cdots+\beta\sqrt{n}<\beta^{2}n$ paths to be added to $\mathcal{P}$, all of which are monotone and, therefore, $\mathcal{P}$ contains at most $\frac{2\pi}{c}\sqrt{n}\cdot\beta^{2}n=\Theta(n^{3/2})$ monotone paths.
We claim that the paths in $\mathcal{P}$ cover the complete geometric graph on $A$. That is, for any pair of points $a,b\in A$, we show that the segment $ab$ appears in at least one path in $\mathcal{P}$. We construct the region $R(a,b)$ analogous to the one in the proof of Theorem 1.1 but with respect to $\varphi$ instead. By Lemma 2.1, its area is at most $4\varphi$. Assume that $R(a,b)$ contains $k$ points of $A$. The disks of radius $r=\frac{\sqrt{2}}{2}\cdot\alpha^{-1}n^{-1/2}$ centered at these $k$ points are disjoint and they cannot exceed the region $R(a,b)$ extended by a margin of $r$, whose area is then at most $4(\varphi+r)$. The total area of the $k$ disks is at most $4(\varphi+r)$, and therefore
@@
$k\pi$ $\leq 8\alpha^{2}n\left(\frac{c}{\sqrt{n}}+\frac{\sqrt{2}}{2\alpha\sqrt{n}}\right)\text{ or }$
$k$ $\leq\frac{8\alpha^{2}c+4\sqrt{2}\cdot\alpha}{\pi}\cdot\sqrt{n}=\beta\sqrt{n}.$
@@
Lower bound. Assume that $n$ is divisible by $3$, i.e., $n=3k$. Partition the $n$ points in our set $P$ into three groups $A,B,C$, of the same size, close to the vertices of a unit equilateral triangle, where each group consists of roughly equidistant points, but the distances are different; see Fig. 2. Let the diameters of the 3 groups of $k$ vertices be
${\rm diam}(C)\ll{\rm diam}(B)\ll{\rm diam}(A)\ll 1.$
The number of (undirected) inter-group edges, namely edges in
@@
is $3k^{2}=n^{2}/3$. Let $\mathcal{P}$ be a covering of $K_{n}[P]$ by monotone paths. We claim that every path $\xi\in\mathcal{P}$ contains at most five edges in $E_{0}$. Consequently, covering all the edges of the tripartite graph on $P=A\cup B\cup C$ requires at least $3k^{2}/5=n^{2}/15$ monotone paths.
@@
Upper bound. Since every 2-edge path is monotone, one can decompose the edge set of $K_{n}[A]$ into $\lceil\frac{1}{2}\binom{n}{2}\rceil$ 2-edge paths. We can improve upon this bound by using $3$-edge paths. Since not every $3$-edge path is monotone, we need special $3$-edge paths that are monotone. A $3$-edge path is a zig-zag path if its two endpoints lie in opposite open halfplanes determined by the middle edge of the path. It is easy to see that any $3$-edge zig-zag path is monotone.
@@
Let $n$ be a large enough positive integer. Then the edge set of $K_{n}$ can be decomposed into subgraphs of the form $K_{6}$, if and only if $n$ satisfies two divisibility conditions, namely 5|(n1)5\ |\ (n-1) and $\binom{6}{2}\ |\ \binom{n}{2}$.
@@
Figure 3: Covering the edge set of the complete graph on six points by 33-edge zig-zag paths.
Our approach is packing as many $K_{6}$ as possible into $K_{n}[A]$, then decomposing each $K_{6}$ into five $3$-edge zig-zag paths, and considering each of the remaining edges as a single path. This way we can guarantee a decomposition of the edge set of $K_{n}[A]$ into $\binom{n}{2}/\binom{6}{2}\times 5+O(n)=n^{2}/6+O(n)$ monotone paths. It may be worth noting that if 5|(n1)5\ |\ (n-1), the second condition simplifies to 3|n(n1)3\ |\ n(n-1).
It suffices to show that for a $6$-point set $A$, the edge set of $K_{6}[A]$ can be decomposed into five $3$-edge zig-zag paths. It is known that there are $16$ order types for such sets, see, e.g., [16]. Since the existence of a zig-zag path decomposition depends only on the order type and not on the specific geometric realization, it suffices to provide zig-zag path decompositions for each type, as we do in Fig. 3. ∎
A similar bipartite version of the lower bound construction with $n=2k$ points and $|E_{0}|=k^{2}$ (undirected) inter-group edges only yields a lower bound of $n^{2}/16$: If $\mathcal{P}$ is a covering of $K_{n}[A]$ by monotone paths, it has the property that every path $\xi\in\mathcal{P}$ contains at most four edges in $E_{0}$. Consequently, covering all the edges of the bipartite graph requires at least $k^{2}/4=n^{2}/16$ monotone paths.